# Converting Transformers into DGNNs Form

### v1

@@ Appendix B Synvolution and Other Attention Mechanisms
We compare Synvolution with common attention and attention-like mechanisms in Table 5. Notice, we do not assume the relation between the length size $N$ and embedded size $D$ for input signal $\mathbf{X}\in\mathbb{R}^{N\times D}$. Sparse attention, which reduces computation by only attending to a subset of tokens, includes [Child et al., 2019, Zaheer et al., 2020]. Following the categorization in [Cao, 2021], linear attention can be classified into Fourier-type and Galerkin-type. Fourier-type attention, which compute query-key pairs first, include [Shen et al., 2021, Cao, 2021, Zhang et al., 2021, Koohpayegani and Pirsiavash, 2024]. Galerkin-type attention, compute key-value pairs first, include [Katharopoulos et al., 2020, Cao, 2021, Koohpayegani and Pirsiavash, 2024]. Kernel attention, which leverages kernel methods to approximate the attention operation, includes [Tsai et al., 2019, Choromanski et al., 2021, Chen et al., 2021b, Xiong et al., 2021]. Synthetic dense attention includes [Tay et al., 2021a]. Chord attention includes [Yu et al., 2022a].
@@
Synvolution Learnable, Full-Rank, and Dense $\mathcal{O}(ND^{2})$ $\mathcal{O}(ND)$
@@ Appendix E Proofs
    Multiplication with $\mathbf{H}^{(l)}_{\text{u}}$ requires $\mathcal{O}(N)$ operations
@@
    Multiplication with $\mathbf{H}^{(l)}_{\text{l}}$ requires $\mathcal{O}(N)$ operations
@@
As shown in Algorithm 1, these multiplications can be implemented efficiently through Hadamard products, the total time complexity is $\mathcal{O}(LN)$. ∎

### v3

@@ Appendix B Synvolution and Other Attention Mechanisms
We compare Synvolution with common attention and attention-like mechanisms in Table 5. Notice, we do not assume the relation between the length size $N$ and embedded size $D$ for input signal $\mathbf{X}\in\mathbb{R}^{N\times D}$. Sparse attention, which reduces computation by only attending to a subset of tokens, includes [Child et al., 2019, Zaheer et al., 2020]. Following the categorization in [Cao, 2021], linear attention can be classified into Fourier-type and Galerkin-type. Fourier-type attention, which compute query-key pairs first, includes [Shen et al., 2021, Cao, 2021, Zhang et al., 2021, Koohpayegani and Pirsiavash, 2024]. Galerkin-type attention, compute key-value pairs first, includes [Katharopoulos et al., 2020, Cao, 2021, Koohpayegani and Pirsiavash, 2024]. Kernel attention, which leverages kernel methods to approximate the attention operation, includes [Tsai et al., 2019, Choromanski et al., 2021, Chen et al., 2021b, Xiong et al., 2021]. Synthetic dense attention includes [Tay et al., 2021a]. Chord attention includes [Yu et al., 2022a].
@@
Synvolution Learnable, Full-Rank, and Dense $\mathcal{O}(ND\log{N}+ND^{2})$ $\mathcal{O}(ND)$
@@ Appendix E Proofs
    Multiplication with $\mathbf{H}^{(l)}_{\text{u}}$ requires $\mathcal{O}(N\log{N})$ operations
@@
    Multiplication with $\mathbf{H}^{(l)}_{\text{l}}$ requires $\mathcal{O}(N\log{N})$ operations
@@
As shown in Algorithm 2, these multiplications can be implemented efficiently through Hadamard products, the total time complexity is $\mathcal{O}(LN\log{N})$. ∎